OpenStudy (anonymous):

Find equations of the lines passing through (-3,5) and having the following characteristics. (a) Slope of 7/16 (b) Parallel to the line 5x-3y=3

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@love_jessika15

OpenStudy (anonymous):

You know y= mx + b you are given the slope.. m= 7/16 so now just plug in a know point with x and y values to find b (the point where the line crosses the y-axis)

OpenStudy (anonymous):

5=(7/16 )(-3)?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

i know the answeris 7x -16y +101 =0 but i dont know how to get it